OpenStudy (anonymous):

simplify the expression (-4p)^2

OpenStudy (lgbasallote):

hint: \[\huge a^{-m} \implies \frac{1}{a^m}\] does that help?

OpenStudy (anonymous):

not sure what am i suppose to do...

OpenStudy (lgbasallote):

what is your a in this case?

OpenStudy (anonymous):

-4

OpenStudy (lgbasallote):

it's actually -4p <--that's your base

OpenStudy (lgbasallote):

and what is m?

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

right. so what is the value of \[\huge \frac{1}{a^m}\]

OpenStudy (anonymous):

so I just rewrite it in that form

OpenStudy (lgbasallote):

yes. remember that a is -4p and m is 2

OpenStudy (anonymous):

oh ok thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (asnaseer):

I don't think this is correct

OpenStudy (asnaseer):

the power here is not negative, your question is:\[(-4p)^2\]

OpenStudy (asnaseer):

which has a positive power

OpenStudy (asnaseer):

@lgbasallote - your example is for negative powers

OpenStudy (lgbasallote):

wait...wasnt that negative a while ago?

OpenStudy (asnaseer):

I don't know? was it @aea2213 ?

OpenStudy (anonymous):

am confused now lol

OpenStudy (asnaseer):

lgbasallote was asking if you had changed the question. did you originally post this question as:\[(-4p)^{-2}\]

OpenStudy (anonymous):

no the exponent is not negative

OpenStudy (asnaseer):

ok, so then:\[(-4p)^2=(-4p)\times(-4p)=?\]

OpenStudy (anonymous):

so its 16

OpenStudy (asnaseer):

what about the terms involving 'p'?

OpenStudy (anonymous):

16p

OpenStudy (asnaseer):

\[p\times p=?\]

OpenStudy (anonymous):

p

OpenStudy (asnaseer):

:/

OpenStudy (asnaseer):

recall your previous question where I stated the laws of indices

OpenStudy (asnaseer):

\[p\times p=p^1\times p^1=?\]

OpenStudy (anonymous):

p^2

OpenStudy (asnaseer):

\[(-4p)^2=(-4p)\times(-4p)=?\]

OpenStudy (anonymous):

16p^2

OpenStudy (asnaseer):

well done! that is correct. :)

OpenStudy (anonymous):

yay lol thanks

OpenStudy (asnaseer):

yw :)